PSNFSS: test most of the ‘35’ PostScript fonts

Sebastian Rahtz June 17th 1994

This is set in in Adobe AvantGarde (pag):

an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

This is set in in Adobe Bookman (pbk):

an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

This is set in in Adobe Courier (pcr):
an italicized fée; a smallcaps fée; a slanted fée;
a bold fée; a bold italized fée; a bold smallcaps fée;
a bold slanted fée;

This is set in in Adobe Helvetica (phv):

an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

This is set in in Adobe NewCenturySchoolbook (pnc): an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

This is set in in Adobe Palatino (ppl):

an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

This is set in in Adobe Times Roman (ptm): an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

This is set in in Adobe ZapfChancery (pzc): an italicized fée; a smallcaps fée; a slanted fée; a bold fée; a bold italized fée; a bold smallcaps fée; a bold slanted fée;

In Helvetica we can try the narrow fonts.^{1} a condensed fée.

1Thanks to Frank Poppe for putting this in.